# Multiple Instance Verification

### v1

@@ 3.1 Three example MI verification tasks
To enable a better understanding of MI verification and its usefulness, we first present detailed information on the three MI verification tasks illustrated in Figure 3. These are tasks for which we were able to obtain publicly available data for our experiments.
@@
To generate the exemplar labels for handwriting verification (Figure 1(a)), we used an extended version of MNIST, the QMNIST dataset [38], which includes the writer IDs of all the handwritten digits in the MNIST dataset. To differentiate our task from the standard MNIST classification task of digit recognition, and to focus on the query class (i.e., the writer ID), we constructed each exemplar using the same digit across the entire exemplar comprising the query and the target bag. To avoid trivial cases that the query is also in the bag, we ensured that all images in the bag are different from the query even if they may be written by the same writer. Note that the “inference” stage in Figure 1(a) depicts the two goals of our study: verification and explanation.
@@
In the signature verification task (Figure 1(b)), a target bag is composed of multiple authentic signatures, usually called “anchor” images. There are three types of query signature:
@@
For all three tasks described above, we were able to collect identifiers of the key instances, either by construction or from raw data. This information is used during testing, to evaluate the explanation quality of the models, but not during training.
3.2 Formal definition of MI verification
To formally state the task of MI verification, we use the following notation. The (unobservable) class label of an instance $x$ is denoted by $y\in c$ where $c=\{1,2,\cdots,L\}$ and $L$ is the number of unobservable classes. Note that $L$ can be $\infty$. The query instance and its unobservable class label are denoted by $x_{Q},y_{Q}$, and the bag of target instances and their unobservable class labels by $\{x_{1},x_{2},\cdots,x_{N}\}$ and $\{y_{1},y_{2},\cdots,y_{N}\}$ respectively. Unless stated otherwise, in this section, we assume the instances, $x_{Q}$, $x_{n}(n=1,\cdots,N)$, have all been embedded in an $F$-dimensional feature space, that is, $\in\mathbb{R}^{F}$.
Definition 1 (MI verification exemplar).
@@ 3.2 Formal definition of MI verification
[…]
It is worth noting that an MI verification task can be converted into a “multiple pairs-of-instances” learning problem, analogous to MIL, by pairing the query instance with each instance in the target bag. We state this formally as a lemma below.
Lemma 1 (Conversion of MI verification exemplars).
An MI verification exemplar in Definition 1 can be represented as a bag of paired instances $\{(x_{Q},x_{1}),(x_{Q},x_{2}),\cdots,(x_{Q},x_{N})\}$, each with an unobservable binary class label $\hat{y}_{n}\in\{0,1\},n=1,2,\cdots,N$ defined as
[…]
The bag label is given by
[…]
The proof is trivially obtained by construction.
3.3 Permutation invariance in MI verification
In MI verification problems, like classic MIL, the order of the instances in a bag should not matter—the function used to score bags of data should be permutation-invariant. Lemma 1 allows us to borrow theory from existing MIL literature, e.g. [17, Theorem 1], to provide a sufficient and necessary condition for the scoring function, $S(X)$, of a bag of instance-pairs to be permutation invariant. We state this formally in Lemma 2, simply replacing an instance $x_{n}$ with a pair of instances $(x_{Q},x_{n})$ in [17, Theorem 1].
Lemma 2 (Condition for permutation invariance).
@@ 3.3 Permutation invariance in MI verification
where $H_{j}(x_{Q})\in\mathbb{R}^{1\times\frac{F}{h}},\alpha_{j}\in\mathbb{R}^{\frac{F}{h}}$ denote the $j^{th}$ segment of the corresponding vectors.
The first line of Equation (3.3) states that the logit of our models is the sum of “sub-logits” from multiple heads. The sub-logit from each head is the attention-score weighted sum of $N$ Siamese-twin similarities computed from $N$ pairs, where each pair is formed by the query (transformed by $H$) and the corresponding instance in the target bag (transformed by $G$)—conforming to the conversion process in Lemma 1. The last line is obtained by simply treating $\sum_{j}^{h}$ as $g(\cdot)$, and everything inside the summation $\sum_{n}^{N}$ as $f(\cdot,\cdot)$.
Although with theoretical merits, the results in this section are not particularly helpful in guiding us to find specific functional forms of $S(X)$. Indeed, even naive employment of existing SOTA MIL attention mechanisms may fit with Lemma 1 and 2. Intuitively, this naive approach may be problematic because these mechanisms do not account for the special role of the query instance. To substantiate this intuition, we need a new framework for interpreting attention scores.
@@ 3.4 A probabilistic interpretation of attention scores
We present a probabilistic interpretation of attention scores based on the observation that they are non-negative and sum to 1. Therefore, they form a probability distribution $Pr(U)$ of a discrete random variable $U\in\{n\mid n=1,2,\cdots,N\}$ that encodes the relevance to the exemplar label of the $N$ instances it contains. This enables the use of two concepts from probability theory. First, the attention-weighted sum of the instances’ feature vectors can now be seen as an expectation, where this expected vector represents the entire bag. Second, the entropy $H(U)$ of $Pr(U)$ can be used to measure the informativeness of a set of attention scores—large entropy corresponds to an even distribution of attention scores, which provides little information about the relevance of individual instances to the bag’s label—and well-known properties from information theory [8] can be applied.
The probability distribution for a particular MI verification exemplar clearly depends on the information in the query instance and the target bag. We assume that the query instance and the instances in the target bag have all been embedded in an $F$-dimensional feature space, yielding $V^{Query}\in\mathbb{R}^{F}$ and $V^{Target}\in\mathbb{R}^{N\times F}$. It can be shown that, under mild conditions, we can strictly reduce uncertainty, and thus strictly increase informativeness regarding the key instances in a target bag, by considering information from both sources, $V^{Target}$ and $V^{Query}$, rather than just one. We state this as Proposition 1, together with Assumption 1, as follows.
Assumption 1 (Uncertainty of conditioning variables’ informativeness).
@@
Assumption 1 states that informativeness of attention scores brought by $V^{Query}$ and $V^{Target}$ should be uncertain, that is, there should be no guarantee that one variable is more informative than another, and vice versa—any scenario is possible with nonzero probability. [^4] Note this probability is well-defined because a conditional entropy is a random variable. This rules out any special condition or constraint on the marginal informativeness of any conditioning variable. In practice, it is a reasonable assumption because it is generally impossible to say with certainty that any given information source is more or less informative than any other.
Proposition 1 (Variables incorporated in the attention scores).
@@
In particular, conditioning on $V^{Target}$ alone can only reduce the ability to identify key instances correctly. This is important because tackling MI verification using naive combinations of Siamese networks and SOTA attention-based MIL models does not incorporate the information from the query into attention. In contrast, our approach uses both the query instance and the target bag to model the attention scores, and Proposition 1 provides a theoretical foundation for it.
@@ 4 Model architectures
We now describe the different modelling approaches to MI verification that we compare in our study, including the two attention functions used in the new approach based on cross-attention pooling (CAP).
The baseline, benchmark, and CAP-based models compared in our study share the same architecture for the first two layers and the final two layers of the neural networks. In the first two layers, we obtain fixed-length feature vectors by embedding the query instance and all instances in the target bag using appropriate feature extractors from the CV or NLP domains, as required by the application at hand, followed by a layer normalization (LayerNorm) [1]. We denote the feature vectors after LayerNorm as $v^{query}\in\mathbb{R}^{1\times C}$ and $v^{target}\in\mathbb{R}^{N\times C}$ respectively, where $N$ is the number of instances in a bag (“bag size”), and $C$ is the number of channels. Note that bag size $N$ varies across exemplars. For the final layer, regardless of architecture, we adopt a binary cross-entropy loss, in which each exemplar’s prediction is $p=sigmoid(sim)$, where $sim$ is a Siamese-twin similarity score:
@@
where $\alpha\in\mathbb{R}^{1\times C}$ is a vector of learnable parameters, and $v^{Q},v^{P}\in\mathbb{R}^{1\times C}$ are the two outputs of the Siamese twins, corresponding to (a) the encoded query instance and (b) the bag-level representation, which is the attention-weighted sum of the encoded target bag instances, respectively (cf. Figure 2). The function $f(v^{P}_{i},v^{Q}_{i})$ measures the similarity between $v^{P}$ and $v^{Q}$. Equation (6) is the same as that used by [20], except that we use the product function, $f(v^{P}_{i},v^{Q}_{i})=v^{P}_{i}v^{Q}_{i}$ instead of $L_{1}$ distance, which yields
@@
4.1 Our approach: Cross Attention Pooling
@@
(c) Multi-head squeeze and co-excitation
Figure 2: Model architecture for “Cross Attention Pooling”, shown in (2(a)), and its key components: (2(b)) multi-head attention scores and (2(c)) squeeze & co-excitation, best viewed in colour.
The main difference between the Siamese networks investigated in this study is how they perform MI pooling of $v^{target}$ to get $v^{P}$. Our new pooling method CAP is motivated by two considerations. First, as implied by Proposition 1, it is strictly better to incorporate $v^{query}$ to model attention scores for the pooling of $v^{target}$. To operationalize this idea, we construct the attention score for each instance in the target bag ($v^{target}$) by attending it to $v^{query}$ as a cross-reference, inspired by the “query-key-value” framework in the transformer decoder [34].
The second consideration comes from the challenge of handling tasks in which all instances in a bag look alike, exemplified by the QMNIST problem in this study. The challenge is how to make key instance(s) stand out among similar-looking instances, through an uneven distribution of their attention scores. We tackle this using two strategies. One is to create new functions that provide unevenly distributed attention scores. Another is to adjust channels of the “value”-vector. The former gives rise to our novel attention functions, namely variance-excited multiplicative attention (VEMA) and distance-based attention (DBA). The latter leads to our multi-head adaptation of “squeeze and co-excitation” [15], a variation of the “squeeze and excitation” network [16].
Figure 2 depicts the entire Siamese network architecture, including our CAP approach highlighted by the dotted rectangle (Figure 2(a)), together with its key components for computing attention scores (Figure 2(b)) and “squeeze and co-excitation” (Figure 2(c)) highlighted by green and blue colours respectively—those components are explained more formally in Section 4.1.2 below. As shown in the figure, there are two outputs of CAP: one is the bag-level representation $v^{P}$, another is the transformed query instance $v^{Q}$. $v^{P}$ is a function of the triad of “query, key, value” denoted by “Q, K, V”, whereas $v^{Q}$ is a function of $v^{query}$ denoted by “T”, in the figure. The functions producing $v^{P}$ and $v^{Q}$ are formally specified by Equation (8) as follows.
4.1.1 Multi-head Cross Attention Pooling
Letting $\mathbf{Q}=v^{query}$, $\mathbf{K}=\mathbf{V}=v^{target}$, $as(\cdot,\cdot)$ be the attention-score function, $LN(\cdot)$ be layer normalization, $MHSCE(\cdot)$ be “multi-head squeeze and co-excitation”, we model the outputs of the Siamese twins as

### v2

@@ 3.1 Three example MIV tasks
To better understand MIV and its usefulness, we first present detailed information on the three MIV tasks illustrated in Figure 4. These are tasks for which we were able to obtain publicly available data for our experiments.
@@
To generate the exemplar labels for handwriting verification (Figure 1(a)), we used an extended version of MNIST, the QMNIST dataset (Yadav and Bottou (2019)), which includes the writer IDs of all the handwritten digits in the MNIST dataset. To differentiate our task from the standard MNIST classification task of digit recognition, and to focus on the query class (i.e., the writer ID), we constructed each exemplar using the same digit across the entire exemplar comprising the query and the target bag. To avoid trivial cases where the query is also in the bag, we ensured that all images in the bag are different from the query even if they may be written by the same writer. Note that the “inference” stage in Figure 1(a) depicts the two goals of our study: verification and key instance detection.
@@
In the signature verification task (Figure 1(b)), a target bag is composed of multiple authentic signatures, usually called “anchor” images. There are three types of query signatures:
@@
For all three tasks described above, we were able to collect identifiers of the key instances, either by construction or from raw data. This information is only used during testing, to evaluate models’ performance in detecting key instances, but not during training.
3.2 Formal definition of MIV
To formally state the task of MIV, we use the following notation. The (unobservable) class label of an instance $x$ is denoted by $y\in c$ where $c=\{1,2,\cdots,L\}$ and $L$ is the number of unobservable classes. Note that $L$ can be $\infty$ in theory. The query instance and its unobservable class label are denoted by $x_{Q},y_{Q}$, and the bag of target instances and their unobservable class labels by $\{x_{1},x_{2},\cdots,x_{N}\}$ and $\{y_{1},y_{2},\cdots,y_{N}\}$ respectively. Unless stated otherwise, in this section, we assume the instances, $x_{Q}$, $x_{n}(n=1,\cdots,N)$, have all been embedded in an $F$-dimensional feature space, that is, $\in\mathbb{R}^{F}$.
Definition 1 (MI verification exemplar)
@@ 3.2 Formal definition of MIV
$Y=\begin{cases}0,&\emph{iff }y\neq y_{Q}\quad\forall y\in\{y_{1},y_{2},\cdots,y_{N}\}\\ 1,&\emph{otherwise.}\end{cases}$
It should be noted that an MIV task can be converted into a “multiple pairs of instances” learning problem, analogous to MIL, by pairing the query instance with each instance in the target bag. We state this formally as a lemma below.
Lemma 2 (Conversion of MIV exemplars)
An MI verification exemplar in Definition 1 can be represented as a bag of instance-pairs: $\{(x_{Q},x_{1}),(x_{Q},x_{2}),\cdots,(x_{Q},x_{N})\}$, each with an unobservable binary class label $\hat{y}_{n}\in\{0,1\},n=1,2,\cdots,N$ defined as
$\hat{y}_{n}=\begin{cases}0,&\emph{iff }y_{n}\neq y_{Q},\quad n=1,2,\cdots,N\\ 1,&\emph{otherwise.}\end{cases}$
The bag-level class label is given by
$Y=\begin{cases}0,&\emph{iff }\sum_{n}^{N}\hat{y}_{n}=0\\ 1,&\emph{otherwise.}\end{cases}$
The proof is obtained trivially by construction.
3.3 Permutation invariance in MIV
In MIV problems, like traditional MIL, the order of the instances in a bag should not matter—the function used to score bags of data should be permutation-invariant. Lemma 2 allows us to borrow theory from existing MIL literature, e.g., (Ilse et al., 2018, Theorem 1), to provide a sufficient and necessary condition for the scoring function, $S(X)$, of a bag of instance pairs to be permutation invariant. We state this formally in Lemma 3, simply replacing an instance $x_{n}$ with a pair of instances $(x_{Q},x_{n})$ in (Ilse et al., 2018, Theorem 1).
Lemma 3 (Condition for permutation invariance)
@@ 3.3 Permutation invariance in MIV
where $H_{j}(x_{Q})\in\mathbb{R}^{1\times\frac{F}{h}},\alpha_{j}\in\mathbb{R}^{\frac{F}{h}}$ denotes the $j^{th}$ segment of the corresponding vectors.
The first line of Equation (3.3) states that the logit of our models is the sum of “sub-logits” from multiple heads. The sub-logit from each head is the attention-score weighted sum of $N$ Siamese-twin similarities computed from $N$ pairs, where each pair is formed by the query (transformed by $H$) and the corresponding instance in the target bag (transformed by GG)—conforming to the conversion process in Lemma 2. The last line is obtained by simply treating $\sum_{j}^{h}$ as $g(\cdot)$, and everything inside the summation $\sum_{n}^{N}$ as $f(\cdot,\cdot)$.
Although with theoretical merits, the results in this section are not particularly helpful in guiding us to find specific functional forms of $S(X)$. Indeed, even naive employment of existing SOTA MIL pooling mechanisms may fit Lemma 2 and 3. However, intuitively, this naive approach may be problematic because these mechanisms do not account for the special role of the query in attention. To substantiate this intuition, we need a new framework for interpreting attention scores.
@@ 3.4 A probabilistic interpretation of attention scores
We present a probabilistic interpretation of attention scores based on the observation that they are non-negative and sum to 1. Therefore, they form a probability distribution $Pr(U)$ of a discrete random variable $U\in\{n\mid n=1,2,\cdots,N\}$ that encodes the relevance to the exemplar label of the $N$ instances it contains. This enables the use of two concepts from probability theory. First, the attention-weighted sum of the instances’ feature vectors can now be seen as an expectation, where this expected vector represents the entire bag. Second, the entropy $H(U)$ of $Pr(U)$ can be used to measure the informativeness of a set of attention scores—large entropy corresponds to an even distribution of attention scores, which provides little information about the relevance of individual instances to the bag’s label—and well-known properties from information theory (Cover and Thomas (2006)) can be applied.
The probability distribution for a particular MIV exemplar clearly depends on the information in the query instance and the target bag. We assume that the query instance and the instances in the target bag have all been embedded in an $F$-dimensional feature space, yielding $V^{Query}\in\mathbb{R}^{F}$ and $V^{Target}\in\mathbb{R}^{N\times F}$. It can be shown that, under mild conditions, we can strictly reduce uncertainty, and thus strictly increase informativeness regarding the key instances in a target bag, by considering information from both sources, $V^{Target}$ and $V^{Query}$, rather than just one. We state this as Proposition 4, together with Assumption 1, as follows.
Assumption 1 (Uncertainty of conditioning variables’ informativeness)
@@
Assumption 1 states that informativeness of attention scores brought by $V^{Query}$ and $V^{Target}$ should be uncertain, that is, there should be no guarantee that one variable is more informative than another, and vice versa—any scenario is possible with nonzero probability. [^5] Note this probability is well-defined because conditional entropy is a random variable. This rules out any special condition or constraint on the marginal informativeness of any conditioning variable. In practice, it is a reasonable assumption because it is generally impossible to say with certainty that any given information source is more or less informative than any other.
Proposition 4 (Variables incorporated in the attention scores)
@@
In particular, conditioning on $V^{Target}$ alone can only reduce the ability to correctly identify key instances. This is important because tackling MIV using naive combinations of Siamese networks and SOTA attention-based MIL models does not incorporate the information from the query into attention. In contrast, our approach uses both the query instance and the target bag to model the attention scores, and Proposition 4 (including Assumption 1) provides a theoretical foundation for it.
3.5 MIV versus related learning problems
To highlight the significance of MIV, Figure 2 schematically depicts the relationship between application scenarios suitable for MIV and those corresponding to two closely related problem sets: (a) standard learning problems involving queries and (b) applications where traditional MIL is appropriate. [^6] We thank an anonymous reviewer for suggesting the inclusion of this content to position the MIV problem.
Problem MIL MIV Problems
characteristics w/ queries
Class label is for: Bag “Bag vs. “Instance vs.
query” pair query” pair
Are queries involved? No Yes Yes
Figure 2: A Venn diagram illustrating relationships between learning problems suitable for multi-instance verification (MIV), those that are suitable for traditional multi-instance learning (MIL), and learning problems that involve queries (e.g., classic verification, metric learning, standard retrieval problems, and so forth). MIV and MIL problems both require training data with bag-level labels, but differ on the existence of a query component. Classic problems involving queries, typically in the form of “instance vs. query (instance)” pairs, require training labels attached to these pairs of instances—i.e., at the instance-level, as opposed to the bag-level.
3.5.1 MIV vs. standard problems involving queries
Importantly, standard learning problems that involve queries, including metric learning, one-shot learning, and standard retrieval problems, are similar to the classic verification problem, generally not involving the concept of “bags of unlabeled instances” and bag-level class labels. Concretely, they all need training data with class labels attached to the “instance vs. query” pairs, where the query itself is also an instance—hence requiring the ground truth of the instance pairs during training. For example, in the standard retrieval problem (likewise in metric or one-shot learning), a “match” can only be defined if a unit of the searched content can be labeled as “matched”(positive) or “not matched”(negative) with the query, requiring such labels of all instance pairs in the training data. The labels may be explicit (by annotation) or implicit (inferred from the searched content’s location, information sources, etc.), as commonly seen in standard “question answering” or “image retrieval” tasks. Even in self-supervised training, labels of instance pairs are still required—for example, a label indicating whether an instance is (a distortion of) the query is typically constructed for self-supervised metric learning.
The availability of ground-truth labels for instance pairs can be viewed as an instance-level label requirement, which is in stark contrast to the bag-level label requirement of MIV—as stated in Lemma 2 (Section 3.2), an MIV exemplar can be represented as a bag of unlabeled instance pairs, with only a bag-level label available. In practical applications, if the instance-level label requirement is met, the MIV formulation is not needed. Therefore, in terms of practical applications, MIV and standard problems involving queries are applicable to mutually exclusive scenarios, and the appropriate formulation should be adopted depending on the available data.
It is also worth noting that even if instance-level labels exist, it may be appropriate to conceptually view the problem at hand as an MIV task with bag-level labels when annotation errors, embedding quality issues, or constraints from feature extractors render individual instance-level labels unreliable.
3.5.2 MIV vs. MIL
Approaches Motivated by what Novelty in relation to
representation of MIV data? traditional MIL
CAP (ours) A bag of (unlabeled) instances Deviation from traditional MIL
plus a query related to the bag due to additional role of query;
(satisfying Assumption 1) New MI pooling algorithm
Baseline A bag containing (unlabeled) New: handling instance-pairs;
model pairs of instances (Lemma 2) Not new: aggregation to bag-level;
Benchmark A bag plus a query independent of No novelty: naive adaptation of
models the bag (violating Assumption 1) existing MIL algorithms;
Table 1: A summary of approaches considered in our experiments, including their conceptual motivations based on different data representations (or assumptions), and their levels of novelty in relation to the existing MIL literature. Both baseline and benchmark models are motivated by representations more akin to that of traditional MIL, so as to leverage existing MIL algorithms, either partially or fully. In contrast, our approach (CAP) deviates from MIL by focusing on the query’s role in inducing bag-level embeddings, in addition to its classic role in verification (as an “anchor”).
Compared to traditional MIL, MIV exhibits additional complexity and is independently significant. The difference between MIV and MIL rests on the existence of a query component in the feature space. The query is subject to transformations (or assumptions), making MIV a multi-faceted problem comprising multiple representations of the data. Table 1 summarizes three different representations of the data in MIV problems (middle column), which correspond to different approaches considered in our study (leftmost column) and exhibit different levels of novelty, or deviation, in relation to standard MIL (rightmost column).
The benchmark models we consider (in Section 4.2) are motivated by the assumption that the query is independent of the target bag—implying zero probability that the query can be more informative than instances of the target bag—violating Assumption 1 in Section 3.4. This representation makes MIV a trivial extension of MIL, and the query only has its classic role in verification as an “anchor”. Consequently, all benchmark models can naively apply existing MIL algorithms to construct bag-level embeddings without considering the query.
The second representation, giving rise to our baseline model (cf. Section 4.2.1), is that from Lemma 2 in Section 3.2, which is based on bags of instance-pairs. The baseline model handles instance-pairs using Siamese networks and processes bags of instance-pairs in an MIL manner. It is partially novel compared to standard MIL approaches because there is no dedicated mechanism in standard MIL to deal with instance pairs.
In contrast to the baseline and benchmark models, our new CAP-based approach is motivated by a representation satisfying Assumption 1, and consistent with Proposition 4. It focuses on the query’s additional role in inducing bag-level embeddings, resulting in the novel MI pooling algorithm, CAP (see Section 4.1). Notably, this perspective on the role of the query, and the resultant pooling mechanism, do not exist in the MIL literature, rendering MIV distinct from MIL.
Considering the approaches in Table 1, our empirical results show that the more a representation deviates from that of MIL, the more empirical success we can obtain in MIV applications, signifying the importance of choosing an approach based on suitable representations or assumptions—in our paper, it should be Assumption 1, Proposition 4, and proper consideration of the query’s role in inducing bag-level embeddings.
@@ 4 Model architectures
We now describe the different modeling approaches compared in our experiments in detail, including the baseline model, the benchmark models, and the new approach based on cross-attention pooling (CAP), instantiated with two novel attention functions we propose.
All models compared in our study share the same architecture for the first two layers and the final two layers of the neural networks. In the first two layers, we obtain fixed-length feature vectors by embedding the query instance and all instances in the target bag using appropriate feature extractors from the CV or NLP domains, as required by the application at hand, followed by a layer normalization (LayerNorm, Ba et al. (2016)). We denote the feature vectors after LayerNorm as $v^{query}\in\mathbb{R}^{1\times C}$ and $v^{target}\in\mathbb{R}^{N\times C}$ respectively, where $N$ is the number of instances in a bag (“bag size”), and $C$ is the number of channels. Note that bag size $N$ varies across exemplars. For the final layer, regardless of architecture, we adopt a binary cross-entropy loss, in which each exemplar’s prediction is $p=sigmoid(sim)$, where simsim is a Siamese-twin similarity score:
@@
where $\alpha\in\mathbb{R}^{1\times C}$ is a vector of learnable parameters, and $v^{Q},v^{P}\in\mathbb{R}^{1\times C}$ are the two outputs of the Siamese twins, corresponding to (a) the encoded query instance and (b) the bag-level representation, which is the attention-weighted sum of the encoded target-bag instances, respectively (cf. Figure 3). The function $f(v^{P}_{i},v^{Q}_{i})$ measures the similarity between $v^{P}$ and $v^{Q}$. Equation (6) is the same as that used by Koch et al. (2015) except that we use the product function, $f(v^{P}_{i},v^{Q}_{i})=v^{P}_{i}v^{Q}_{i}$ instead of $L_{1}$ distance, which yields
@@
4.1 Our approach: Cross-Attention Pooling (CAP)
@@
(c) Multi-head co-excitation
Figure 3: Model architecture for “Cross-Attention Pooling”, shown in (3(a)), and its key components: (3(b)) multi-head attention scores and (3(c)) multi-head co-excitation, best viewed in colour. In (3(a)), Q, K, V, T denote the vectors of “query”, “key”, “value” and “transformed query instance” (to be transformed into $v^{Q}$) , respectively. In (3(b)), $diag(\delta)$ denotes a $(D\times D)$ diagonal-matrix based on $\delta$ with the function $I_{j}(\cdot)$ defined by Equation (10), Eq. 9 and Eq.11 denote Equations (9) and (11) respectively. In (3(c)), $\odot$ denotes element-wise multiplication, and the function $MHCE(\cdot)$ is defined by Equation (12).
The main difference between the algorithms investigated in this study is how they conduct MI-pooling of $v^{target}$ to get $v^{P}$. Our new pooling method CAP is motivated by two considerations. First, as implied by Proposition 4, it is strictly better to incorporate $v^{query}$ to model attention scores for the pooling of $v^{target}$. To operationalize this idea, we construct the attention score for each instance in the target bag ($v^{target}$) by attending it to $v^{query}$ as a cross-reference, inspired by the “query-key-value” framework in the transformer decoder (Vaswani et al. (2017)).
The second consideration comes from the challenge of handling tasks in which all instances in a bag look alike, exemplified by the QMNIST task in our study. The challenge is how to make key instance(s) stand out among similarly looking instances, through an uneven distribution of their attention scores. We tackle this using two strategies. One is to create new functions that provide unevenly distributed attention scores. Another is to excite channels of the “value”-vector. The former gives rise to our novel attention functions, namely variance-excited multiplicative attention (VEMA) and distance-based attention (DBA). The latter leads to our multi-head adaptation of “co-excitation” (Hsieh et al. (2019)), a variant of the “squeeze and excitation” network (Hu et al. (2018)).
Figure 3 depicts the entire Siamese network architecture, including our CAP approach highlighted by the dotted rectangle (Figure 3(a)), together with its key components for computing attention scores (Figure 3(b)) and “co-excitation” (Figure 3(c)) highlighted by green and blue colours respectively—those components are explained more formally in Section 4.1.2 below. As shown in the figure, there are two outputs of CAP: one is the bag-level representation $v^{P}$, another is the transformed query instance $v^{Q}$. $v^{P}$ is a function of the triad of “query, key, value” denoted by “Q, K, V”, whereas $v^{Q}$ is a function of $v^{query}$ denoted by “T” in the figure. The functions producing $v^{P}$ and $v^{Q}$ are formally specified by Equation (8).
4.1.1 Multi-head Cross-Attention Pooling
Letting $\mathbf{Q}=v^{query}$, $\mathbf{K}=\mathbf{V}=v^{target}$, $as(\cdot,\cdot)$ be the attention-score function, $LN(\cdot)$ be layer normalization, $MHCE(\cdot)$ be “multi-head co-excitation”, we model the outputs of the Siamese twins as